Suppose the firm faces a price of $38, an average variable cost of $29, and has an average fixed cost of $5. In the​ short-run, the firm

A. will earn an economic profit

B. will be unable to determine what to do

c. will just cover cost

D. none of the above
Answer #1

ATC = AVC+AFC = 29 + 5 = 34

P = 38

As P > ATC,

Firm will earn profit

Therefore correct answer is option A